# Feedforward suppression of readout-induced faults in quantum error correction

### v1

@@ Abstract
The presented approach therefore broadens the palette of cost-benefit options with which to improve decoder performance and better suppress effects like readout error and leakage against modest costs including some hardware modifications to enable feedforward, a possible feedforward delay and impact of the additional gate insertion (e.g., idling errors of data qubits and the X gate error). Beyond the directly reduced logical error, the reduction of decoding time could have significant implications, allowing for real-time decoding of larger codes and employing larger check matrices encompassing more faults, eventually further reducing the logical error rates.
Acknowledgements
@@
We thank Luke Govia, Maika Takita, David McKay, Ted Yoder, Tomas Jochym-O’Connor and Lev Bishop for very helpful feedback. Research by H.L. and L.S. was partially sponsored by the Army Research Office and was accomplished under Grant Number W911NF-21-1-0002. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
Appendix A Evolution of the population of leaked qubits
@@
The state in which a check qubit would be measured is modeled as being in one of three levels, the ground state (0), the excited state (1), or a leaked state (2, denoting any higher-energy quantum level). The probability distribution over the three possible states of each check qubit in cycle $i$ is defined by a column vector $\nu_{i}$, with the entries ordered by the probabilities of the states 0, 1 and 2.
@@
where we have defined for brevity $p_{l}\equiv p_{\rm leak},$ and $p_{s}\equiv p_{\rm seep}$. Here, $\delta$ is 0 without the flip protocol, which describes a scenario where check qubits have identical probabilities to be measured in 0 or 1 (due to the Pauli frame being random and ignoring biased readout errors), and qubits leak from (to) the level 1 with the leakage (seepage) probability. With $\delta\neq 0$, the model describes an increased qubit probability to be measured in the state 0, due to the flip protocol, with the exact magnitude of the phenomenological constant $\delta$ depending on other various error parameters of the model and decoding.
@@
Figure 5: The evolution of the mean number of leaked check qubits across the cycles of one syndrome measurement circuit, simulated using the state-dependent Pauli model as described above, with error rates as in Eqs. (2)-(4) and $p_{\rm back}=0.025$. (a), (b) The ground state population for Z and X checks. (c), (d) The leaked state population, for Z and X checks. Starting from an initial state with no leaked qubits, the number of leaked qubits grows and approaches a steady state. The mean population in the leaked state just before readout confirms the reduction of leakage by the flip protocol.
In Fig. 5 we show the evolution of the mean number of leaked check qubits across the cycles of one syndrome measurement circuit, simulated using the state-dependent Pauli model. $\delta$ can be extracted numerically and is found to be roughly $\delta\approx 0.43$, giving a ground state population close to 0.93 and an almost one order of magnitude reduction in the leaked population of the checks. The rate of approach to the steady state is in this regime approximately $p_{\rm seep}$. By initializing the state of the decoding simulation to a population of leaked qubits corresponding to the approximate steady state given above, we find that this is a very good self-consistent approximation, and the qubits remain in a state with a very closed ground-state and leaked population. All simulations in this work are performed starting with a random qubit population taken from the leaked state probability.
@@
Figure 7: The dependence of decoding characteristics on preparation errors occurring with the probability $p_{\rm corr}$ conditioned on a readout error having occurred in the preceding readout step. The error parameters are as in Eqs. (2)-(3) and $p_{\rm id,l}=0.005$ and the panels (a), (b) and (d) as in Fig. 6, while panel (c) shows the mean population of qubits in the ground state before the readout steps. A significant increase in decoding errors and time is observed for even small residual preparation errors. See the text for a discussion.
@@
Appendix B Preparation errors correlated with a readout error
Our state-dependent simulations allow us to probe preparation errors correlated to a readout error, modeled as occurring with some probability $p_{\rm corr}$, conditioned on a readout error having occurred during the immediately preceding readout gate. With a high residual preparation error (of order 1 after a readout error) a dramatic increase in the logical error rates is observed. Although this would be the naive scenario for the mid-circuit measurement and reset protocol known as conditional reset (in which the reset is completely based on the readout result), a mechanism for the reduction of this error is likely to be in place.

### v3

@@ Abstract
The presented approach therefore broadens the palette of cost-benefit options with which to improve decoder performance and better suppress effects like readout error and leakage against modest costs that might include some hardware modifications to enable feedforward, a possible feedforward delay and impact of the additional gate insertion (e.g., idling errors of data qubits and the X gate error). Beyond the directly reduced logical error, the reduction of decoding time could have significant implications, allowing for real-time decoding of larger codes and employing larger check matrices encompassing more faults, eventually further reducing the logical error rates.
@@
Acknowledgements. We thank Luke Govia, Maika Takita, David McKay, Ted Yoder, Tomas Jochym-O’Connor, Michael Beverland, Andrew Cross, Lev Bishop and Doug McClure for very helpful feedback. Research by H.L. and L.S. was partially sponsored by the Army Research Office and was accomplished under Grant Number W911NF-21-1-0002. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
Elements of this work are included in a patent application filed by the International Business Machines Corporation with the US Patent and Trademark Office.
Parameters and their value in the main figures
Parameter Fig. 2 Fig. 3 Fig. 4
$p_{\rm read}$ 0.02 0.02 0.02
$p_{\rm bias}$ 0.005 0.005 0.005
$p_{\rm prep}$ 0.001 0.001 0.001
$p_{\rm id,\,s}$ 0.001 0.001 0.001
$p_{\rm cnot}$ 0.001 0.001 0.001
$p_{\rm id,\,l}$ 0.005 / 0.006 0.005 / 0.006 0.005 - 0.01
$p_{\rm leak}$ 0.002 0.0005 - 0.003 0.002
$p_{\rm seep}$ 0.2 0.33 0.2 - 0.7
$p_{\rm flip}$ 0.0002 0.0002 0.0002
$p_{\rm back}$ 0 - 0.05 0.05 0.025
Table 1: A summary of all the parameters used in the different simulations presented in Figs. 2-4.
Appendix A Optimizing the readout parameters
The adaptations of the readout parameters according to the expected measured state can be optimized in some cases, e.g., when there are competing mechanisms of readout-induced faults, or even in the simpler case of biased readout error. To give a concrete example that is simple to analyze, we can focus on readout errors alone and consider a favorable modification of the discrimination. It can happen that for the “standard” calibration of the readout pulse, leakage is stronger from one qubit state while the readout error is larger for the other state. In such a case, the readout gate can be tailored to each state such that the readout errors and induced faults will be reduced (with some tradeoff between different types of such faults). Each adaptation will have some effect and the overall error rates can be optimized. This approach might be beneficial even if the “standard” errors were symmetric.
For a rudimentary analysis, we assume a standard calibration of the readout resulted for each qubit in a minimal mean readout error $p_{\rm read}$ and some given value of $p_{\rm bias}$ that depends on the qubit and measurement properties. A simple adaptation of the readout would be to set a different threshold of discrimination, decreasing $p(1|0)$ while paying with $p(0|1)$. Defining the prior probability $p_{0}$ of a check qubit to be in the 0 state just before readout at the end of a syndrome cycle, the results of App. C show that with the flip protocol we get $p_{0}\approx 0.93$ (for the parameters described there). Then it is clear that by minimizing $p_{0}\times p(1|0)+p_{1}\times p(0|1)$ in the setting of the QEC syndrome cycle, the overall readout error rate can be reduced. Since $p_{0}$ depends on all errors during the syndrome cycle, the best optimization would require an experiment applying the protocol to known states and adapting the measurement gates such that a minimal overall error (or decoding time) is observed.
Appendix B Preparation errors correlated with a readout error
Figure 5: The dependence of decoding characteristics on preparation errors occurring with the probability $p_{\rm corr}$ conditioned on a readout error having occurred in the preceding readout step. The error parameters are as in Eqs. (9)-(10) and $p_{\rm id,l}=0.005$. (a) The mean number of BP iterations per decoding window. (b) The fraction of BP algorithm runs that converged. (c) The mean population of check qubits in the ground state just prior to the syndrome cycles’ readout step. (d) The logical error rate per cycle. A significant increase in decoding errors is observed for even small residual preparation errors. See the text for a discussion.
Our state-dependent simulations allow us to probe preparation errors correlated to a readout error, modeled as occurring with some probability $p_{\rm corr}$ conditioned on a readout error having occurred in the immediately preceding readout gate. With a high residual preparation error (of order 1 after a readout error) a dramatic increase in the logical error rates is observed. Although this would be the naive scenario for the mid-circuit measurement and reset protocol known as conditional reset (in which the reset is directly based on the readout result), steps for the reduction of this error can be anticipated.
Appendix C Evolution of the population of leaked qubits
@@
The state in which a check qubit would be measured is modeled as being in one of three levels, the ground state (0), the excited state (1), or a leaked state (2, denoting any higher-energy quantum level). The probability distribution over the three possible states of each check qubit in cycle $i$ is defined by a column vector $\nu_{i}$, with the entries in the order of the states 0, 1 and 2.
@@
where we have defined for brevity $p_{l}\equiv p_{\rm leak},$ and $p_{s}\equiv p_{\rm seep}$. Here, $\delta$ is 0 without the flip protocol, which describes a scenario where check qubits have identical probabilities to be measured in 0 or 1 (due to the Pauli frame being random and ignoring biased readout errors), and qubits leak from (seep to) the level 1 with the leakage (seepage) probability. With $\delta\neq 0$, the model describes an increased qubit probability to be measured in the state 0, due to the flip protocol, with the exact magnitude of the phenomenological constant $\delta$ depending on other various error parameters of the model and decoding.
@@
Figure 6: The evolution of the mean number of leaked check qubits across the cycles of one syndrome measurement circuit, simulated using the state-dependent Pauli model as described above, with error rates as in Eqs. (9)-(11) and $p_{\rm back}=0.025$. For the purpose of this figure, no qubits were initialized in a leaked state. (a), (b) The ground state population for Z and X checks. (c), (d) The leaked state population, for Z and X checks. Starting from an initial state with no leaked qubits, the number of leaked qubits grows and approaches a steady state. The mean population in the leaked state just before readout confirms the reduction of leakage by the flip protocol.
@@
In Fig. 6 we show the evolution of the mean number of leaked check qubits across the cycles of one syndrome measurement circuit, simulated using the state-dependent Pauli model, when no check qubits are initially leaked. The error rates are set to
the state-dependent readout parameters are,
$p_{\rm read}=0.02,\quad p_{\rm bias}=0.005,$ (10)
and we have also
$p_{\rm id,\,l}=0.005,\quad p_{\rm leak}=0.002,\quad p_{\rm seep}=0.2.$ (11)
$\delta$ can be extracted numerically and is found to be roughly $\delta\approx 0.43$, giving a ground state population close to 0.93 and an almost one order of magnitude reduction in the leaked population of the checks. The rate of approach to the steady state is in this regime approximately $p_{\rm seep}$. By initializing the state of the decoding simulation to a population of leaked qubits corresponding to the approximate steady state given above, we find that this is a very good self-consistent approximation, and the evolving state maintains a very similar population of the different levels (ground, excited and leaked). All simulations presented in the main text of this work are performed starting with a random qubit population with leakage.